# Chemical Reaction Neural Networks for Fitting Accelerated Rate Calorimetry Data bBeonD Srl - Corso Castelfidardo, 30A, 10129 Torino, Italy)

### v1

@@ Abstract
As the demand for lithium-ion batteries rapidly increases there is a need to design these cells in a safe manner to mitigate thermal runaway. Thermal runaway in batteries leads to an uncontrollable temperature rise and potentially fires, which is a major safety concern. Typically, when modelling the chemical kinetics of thermal runaway calorimetry data ( e.g. Accelerated Rate Calorimetry (ARC)) is needed to determine the temperature-driven decomposition kinetics. Conventional methods of fitting Arrhenius Ordinary Differential Equation (ODE) thermal runaway models to Accelerated Rate Calorimetry (ARC) data make several assumptions that reduce the fidelity and generalizability of the obtained model. In this paper, Chemical Reaction Neural Networks (CRNNs) are trained to fit the kinetic parameters of N-equation Arrhenius ODEs to ARC data obtained from a Molicel 21700 P45B. The models are found to be better approximations of the experimental data. The flexibility of the method is demonstrated by experimenting with two-equation and four-equation models. Thermal runaway simulations are conducted in 3D using the obtained kinetic parameters, showing the applicability of the obtained thermal runaway models to large-scale simulations.
Keywords: Thermal Runaway, Accelerated Rate Calorimetry (ARC), Artificial Intelligence (AI), Chemical Reaction Neural Networks (CRNN), Neural ODE
@@ 1 Introduction
Typically, when modeling thermal runaway, chemical decomposition kinetics are represented by multiple Arrhenius Ordinary Differential Equations (ODEs)[5, 2, 6]. These reaction rates, combined with the reaction enthalpy, allow for the prediction of chemical heat generation under abuse scenarios. For each of the Arrhenius ODEs, the kinetic parameters and heat of reaction must be determined. In the literature, a popular model is that proposed by Hatchard et al. [5]. This model utilized calorimetry-derived data to model thermal runaway in an 18650 Lithium Cobalt Oxide (LCO) cell for the anode, cathode, and SEI decomposition reactions. The model utilized a lumped approach, that is treating the jellyroll as a single temperature, and demonstrated excellent agreement with experimental oven tests. This model has been the basis of a number of subsequent models and has been extended to simulate other cell types, 3D domains (i.e. not lumped), additional decomposition reactions (e.g., electrolyte breakdown and short circuit events), and different cathode chemistries[6, 7, 8, 9, 10]. In addition to these Hatchard-derived models, parameter estimation for the generic Arrhenius ODE have been proposed in the literature. The model by Shelkea et al. [11], which is similar to [5] and [6], but with the addition of a mix term to match the experimental accelerated rate calorimetry (ARC) data, is one example. Ping et al. [12] demonstrated a five-stage model also based on ARC data. The kinetic and reaction heat parameters were determined through a combination of parameter fitting and literature data. More details regarding this method are provided later. Feng et al. [13] proposed an elaborate mathematical model based on literature (e.g. [5]) and some novel formulations, demonstrating excellent agreement with the ARC experimental data. Their model utilized parameters from the literature and adjusted the parameters to fit experimental data; limited details of the fitting methods were provided. Chen et al. [14] introduced an ARC-derived two-stage model. The advantage of this model is it requires fewer parameters than the four-stage NREL model. Their approach demonstrated good agreement with the ARC experimental data. In general, the models currently in literature can intricately model the physics of thermal runaway, but parameter identification often relies on literature data or laborious and not necessarily robust identification methods.
@@
Parameter estimation for thermal runaway modelling remains a major challenge and typically relies on calorimetry data to measure chemical heat release. In literature, there are several types of calorimetry data used, each with their own advantages and disadvantages. Some examples include: Accelerated Rate Calorimetry (ARC)[15, 16, 17, 13, 11, 14]; Fractional Thermal Runaway Calorimetry (FTRC) developed by NASA [18]; and Differential Scanning Calorimetry (DSC) [19, 2, 20]. FTRC can accurately measure heat generation and gas release during thermal runaway, offering the advantage of shorter experimental time frames compared to other calorimetry-based approaches. To the authors’ knowledge, no fitting has been done on this data yet, but it represents a promising method. DSC is a thermal analysis technique that measures heat flow into or out of a sample (e.g. cathode material) as it undergoes a controlled temperature ramp. This heat flow can be used to characterize the heat generation of individual battery components, separately. Finally, ARC is one of the most common ways to characterize battery thermal runaway response for a single cell. In this experiment, the intact cell is first placed inside a thermally isolated cell holder in a sealed calorimeter chamber. The test initiates in a heat-wait-seek mode, where a heater heats the calorimeter incrementally, then pauses to allow thermal transients to decay and equilibrium conditions to be reached. If any increase in temperature rate surpasses a preset threshold value (indicating exothermic reactions), the calorimeter switches to exotherm mode. In this mode, the calorimeter tracks the temperature rise and maintains zero temperature differential between the cell surface and the calorimeter walls, ensuring adiabatic conditions. The final phase involves either full thermal runaway and subsequent cooling or simply cooling if the temperature rate falls below the preset threshold value.
@@
This work proposes a novel method to obtain accurate fits for Arrhenius equation-based thermal runaway models from experimental Accelerated Rate Calorimetry (ARC) data. CRNNs are utilized to learn the required parameters. A staging and linearization-based fitting method is demonstrated on the data, showing its shortcomings in creating accurate models. The obtained parameters are then tuned via training a CRNN, resulting in learned models that better represent the data. The experiment is conducted on two and four-stage models, showing the flexibility of the method. The tuned parameters are tested in 3D simulations of ARC and oven tests, demonstrating improved workflow and performance compared with more traditional methods cited in the literature.
@@ 3.2 Linearized Arrhenius kinetics
Utilizing the staged experimental data, the slope and intercept of the v/s plot return the value of and respectively.
@@ 4.1 Two-stage model
The description of the staging and linearization is identical to Section 3.2 and is omitted here for brevity. The temperatures used for the staging, as shown in Figure (2), were T, T, and T. Figure (5) shows the integrated solution for parameters initially returned from the linear fit (Figs. (5(a)),(5(c))), and for parameters after training a CRNN (Figs. (5(b)), (5(d)) ). Figure (5(a)) shows the temperature increase with time compared to the experimental data, and it is observed that there is a significant difference in the profile of temperature evolution. An overprediction of temperature is observed until the system is close to thermal runaway, where it turns to an underprediction and the thermal runaway time is not correctly predicted. Figure (5(c)) shows the heat rate plotted against temperature, and a significant deviation of the prediction is observed with respect to the experimental data.
@@ 4.2 Four-stage model
The parameters obtained for the linear fit were used to initialize the CRNN training process. Figure (6(b)) shows the temperature evolution with time. The fit qualitatively shows good agreement with the experimental data, both in the self-heating phase and during the rapid temperature rise of thermal runaway. Additionally, the fitted data shows a better prediction of the time the system went into thermal runaway and the overall peak temperature. As an additional observation, the obtained temperature profile for the four-stage fit (Figure 6(b)) adheres better to the experimental data compared with the two-stage fit (Figure 5(b)). This is attributed to the four stages giving the model more ability to accurately represent the data.
@@
Furthermore, as noted in Table 4, the linear fit for the final stage in a four-stage model can be so poor that some parameters from stage 3 need to be used as initial guesses. Sun et al. [31] and Feng et al. [37] instead modeled the heat release of the final stage using a cumulative enthalpy method, which measures the total energy released by the last stage of the process and reuses the same function in future simulations. While this method is simple to implement, it does not generalize well to simulations outside of the ARC test because it does not attempt to model the dynamics of the stage. The CRNN method, however, is able to automatically find parameters governing the dynamics of stage 4, removing the need for approximations such as the cumulative enthalpy method.
@@ 5 Conclusion
This work explored using Chemical Reaction Neural Networks (CRNNs) to perform kinetic parameter fitting of multi-stage thermal runaway models governed by the Arrhenius equations to ARC data. The issues with using linearization to obtain the kinetic parameters were discussed and demonstrated, and the linear fit results were improved using a CRNN for both two-stage and four-stage models. The obtained parameters showed much better agreement with experimental data. The four-stage fit was also used to simulate ARC and oven tests in 3D, with results showing agreement with experiments. This further demonstrated the applicability of the derived thermal runaway models to simulations involving batteries or battery packs, where thermal runaway is a major safety concern and hence is a crucial design factor. The proposed method is flexible and can generalize to several stages, cell types, and stoichiometries, representing a step toward enhanced battery pack safety through simulation-driven design.

### v2

# Chemical Reaction Neural Networks for Fitting Accelerating Rate Calorimetry Data bBeonD Srl - Corso Castelfidardo, 30A, 10129 Torino, Italy)

@@ Abstract
As the demand for lithium-ion batteries rapidly increases there is a need to design these cells in a safe manner to mitigate thermal runaway. Thermal runaway in batteries leads to an uncontrollable temperature rise and potentially battery fires, which is a major safety concern. Typically, when modelling the chemical kinetics of thermal runaway calorimetry data ( e.g. Accelerating Rate Calorimetry (ARC)) is needed to determine the temperature-driven decomposition kinetics. Conventional methods of fitting Arrhenius Ordinary Differential Equation (ODE) thermal runaway models to ARC data make several assumptions that reduce the fidelity and generalizability of the obtained model. In this paper, Chemical Reaction Neural Networks (CRNNs) are trained to fit the kinetic parameters of N-equation Arrhenius ODEs to ARC data obtained from a Molicel 21700 P45B. The models are found to be better approximations of the experimental data. The flexibility of the method is demonstrated by experimenting with two-equation and four-equation models. Thermal runaway simulations are conducted in 3D using the obtained kinetic parameters, showing the applicability of the obtained thermal runaway models to large-scale simulations.
Keywords: Thermal Runaway, Accelerating Rate Calorimetry (ARC), Artificial Intelligence (AI), Chemical Reaction Neural Networks (CRNN), Neural ODE
@@ 1 Introduction
Typically, when modeling thermal runaway, chemical decomposition kinetics are represented by multiple Arrhenius Ordinary Differential Equations (ODEs)[5, 2, 6]. These reaction rates, combined with the reaction enthalpy, allow for the prediction of chemical heat generation under abuse scenarios. For each of the Arrhenius ODEs, the kinetic parameters and heat of reaction must be determined. In the literature, a popular model is that proposed by Hatchard et al. [5]. This model utilized calorimetry-derived data to model thermal runaway in an 18650 Lithium Cobalt Oxide (LCO) cell for the anode, cathode, and SEI decomposition reactions. The model utilized a lumped approach, that is treating the jellyroll as a single temperature, and demonstrated excellent agreement with experimental oven tests. This model has been the basis of a number of subsequent models and has been extended to simulate other cell types, 3D domains (i.e. not lumped), additional decomposition reactions (e.g., electrolyte breakdown and short circuit events), and different cathode chemistries[6, 7, 8, 9, 10]. In addition to these Hatchard-derived models, parameter estimation for the generic Arrhenius ODE have been proposed in the literature. The model by Shelkea et al. [11], which is similar to [5] and [6], but with the addition of a mix term to match the experimental accelerating rate calorimetry (ARC) data, is one example. Ping et al. [12] demonstrated a five-stage model also based on ARC data. The kinetic and reaction heat parameters were determined through a combination of parameter fitting and literature data. More details regarding this method are provided later. Feng et al. [13] proposed an elaborate mathematical model based on literature (e.g. [5]) and some novel formulations, demonstrating excellent agreement with the ARC experimental data. Their model utilized parameters from the literature and adjusted the parameters to fit experimental data; limited details of the fitting methods were provided. Chen et al. [14] introduced an ARC-derived two-stage model. The advantage of this model is it requires fewer parameters than the four-stage NREL model. Their approach demonstrated good agreement with the ARC experimental data. In general, the models currently in literature can intricately model the physics of thermal runaway, but parameter identification often relies on literature data or laborious and not necessarily robust identification methods.
@@
Parameter estimation for thermal runaway modelling remains a major challenge and typically relies on calorimetry data to measure chemical heat release. In literature, there are several types of calorimetry data used, each with their own advantages and disadvantages. Some examples include: ARC [15, 16, 17, 13, 11, 14]; Fractional Thermal Runaway Calorimetry (FTRC) developed by NASA [18]; and Differential Scanning Calorimetry (DSC) [19, 2, 20]. FTRC can accurately measure heat generation and gas release during thermal runaway, offering the advantage of shorter experimental time frames compared to other calorimetry-based approaches. To the authors’ knowledge, no fitting has been done on this data yet, but it represents a promising method. DSC is a thermal analysis technique that measures heat flow into or out of a sample (e.g. cathode material) as it undergoes a controlled temperature ramp. This heat flow can be used to characterize the heat generation of individual battery components, separately. Finally, ARC is one of the most common ways to characterize battery thermal runaway response for a single cell. In this experiment, the intact cell is first placed inside a thermally isolated cell holder in a sealed calorimeter chamber. The test initiates in a heat-wait-seek mode, where a heater heats the calorimeter incrementally, then pauses to allow thermal transients to decay and equilibrium conditions to be reached. If any increase in temperature rate surpasses a preset threshold value (indicating exothermic reactions), the calorimeter switches to exotherm mode. In this mode, the calorimeter tracks the temperature rise and maintains zero temperature differential between the cell surface and the calorimeter walls, ensuring adiabatic conditions. The final phase involves either full thermal runaway and subsequent cooling or simply cooling if the temperature rate falls below the preset threshold value.
@@
This work proposes a novel method to obtain accurate fits for Arrhenius equation-based thermal runaway models from experimental ARC data. CRNNs are utilized to learn the required parameters. A staging and linearization-based fitting method is demonstrated on the data, showing its shortcomings in creating accurate models. The obtained parameters are then tuned via training a CRNN, resulting in learned models that better represent the data. The experiment is conducted on two and four-stage models, showing the flexibility of the method. The tuned parameters are tested in 3D simulations of ARC and oven tests, demonstrating improved workflow and performance compared with more traditional methods cited in the literature.
@@ 3.2 Linearized Arrhenius kinetics
Utilizing the staged experimental data, the slope and intercept of the v/s plot return the value of and , respectively.
@@ 4.1 Two-stage model
The description of the staging and linearization is identical to Section 3.2 and is omitted here for brevity. The temperature values used for the staging, as shown in Figure (2), were T, T, and T. Figure (5) shows the integrated solution for parameters initially returned from the linear fit (Figs. (5(a)),(5(c))), and for parameters after training a CRNN (Figs. (5(b)), (5(d)) ). Figure (5(a)) shows the temperature increase with time compared to the experimental data, and it is observed that there is a significant difference in the profile of temperature evolution. An overprediction of temperature is observed until the system is close to thermal runaway, where it turns to an underprediction and the thermal runaway time is not correctly predicted. Figure (5(c)) shows the heat rate plotted against temperature, and a significant deviation of the prediction is observed with respect to the experimental data.
@@ 4.2 Four-stage model
The parameters obtained for the linear fit were used to initialize the CRNN training process. Figure (6(b)) shows the temperature evolution with time. The fit qualitatively shows good agreement with the experimental data, both in the self-heating phase and during the rapid temperature rise of thermal runaway. Additionally, the fitted data shows a better prediction of the time the system went into thermal runaway and the overall peak temperature. As an additional observation, the obtained temperature profile for the four-stage fit (Figure 6(b)) compares better to the experimental data than with the two-stage fit (Figure 5(b)). This is attributed to the four stages giving the model more ability to accurately represent the data.
@@
Furthermore, as noted in Table 4, the linear fit for the final stage in a four-stage model can be so poor that some parameters from stage 3 need to be used as initial guesses. Sun et al. [31] and Feng et al. [37] instead modeled the heat release of the final stage using a cumulative enthalpy method, which measures the total energy released by the last stage of the process and reuses the same function in future simulations. While this method is simple to implement, it does not generalize well to simulations outside of the ARC test as it does not attempt to model the dynamics of the stage. The CRNN method, however, is able to automatically find parameters governing the dynamics of stage 4, removing the need for approximations such as the cumulative enthalpy method.
@@ 5 Conclusion
This work explored using Chemical Reaction Neural Networks (CRNNs) to perform kinetic parameter fitting of multi-stage thermal runaway models governed by the Arrhenius equations to ARC data. The issues with using linearization to obtain the kinetic parameters were discussed and demonstrated, and the linear fit results were improved using a CRNN for both two-stage and four-stage models. The parameters obtained showed much better agreement with the experimental data. The four-stage fit was also used to simulate ARC and oven tests in 3D, with results showing agreement with experiments. This further demonstrated the applicability of the derived thermal runaway models for simulations involving batteries or battery packs, where thermal runaway is a major safety concern and hence is a crucial design factor. The proposed method is flexible and can generalize to several stages, cell types, and stoichiometries, representing a step toward enhanced battery pack safety through simulation-driven design.